OpenStudy (macabbey101):

What is the solution to the equation? 4x – 6 = –3x + 8

OpenStudy (anonymous):

7X = 14. X = 2

OpenStudy (anonymous):

3x + 4x - 6 = 8 7x - 6 = 8 7x = 8 + 6 7x = 14 x = 2

OpenStudy (macabbey101):

thank you so much!!!:)